I dont have an exact question but if I have an iceberg that is half submerged, how would I go about finding the force required to fully submerge it? can you please explain this using the sum of forces and boyancy ?
Answer #1

If the ice berg is half submerged that means the weight of it is balanced by the weight of liquid displaced by its half volume lets say 50cm3(taking total as 100cm3) which is the boyant force acting on it. To submerge the whole of it we need to apply force equal to weight of liquid displaced by other half of it.

m = 50cm3 * 1g/cm3 = 50 g

Weight = mg = 0.050*9.8 = 0.49 N